### v2

# kooplearn: A Scikit-Learn Compatible Library of Algorithms for Evolution Operator Learning

Abstract
kooplearn is a machine-learning library that implements linear, kernel, and deep-learning estimators of dynamical operators and their spectral decompositions. kooplearn can model both discrete-time evolution operators (Koopman/Transfer) and continuous-time infinitesimal generators. By learning these operators, users can analyze dynamical systems via spectral methods, derive data-driven reduced-order models, and forecast future states and observables. kooplearn’s interface is compliant with the scikit-learn API (Pedregosa et al., 2011), facilitating its integration into existing machine learning and data science workflows. Additionally, kooplearn includes curated benchmark datasets to support experimentation, reproducibility, and the fair comparison of learning algorithms. The software is available at https://github.com/Machine-Learning-Dynamical-Systems/kooplearn.
Statement of Need
From fluid flows down to atomistic motions, dynamical systems permeate every scientific discipline. Among the data-driven frameworks for modeling dynamical systems, evolution operator learning (Kostic et al., 2022) is both general and principled, and is especially well suited for interpretability (Schütte et al., 2001; Mezić, 2005) and dimensionality reduction (Klus et al., 2018). An evolution operator $\mathsf{E}$ characterizes dynamical systems, either stochastic $x_{t+1}\sim p(\cdot|x_{t})$, or deterministic $x_{t+1}\sim\delta(\cdot-F(x_{t}))$, as follows: for every function $f$ of the state of the system, $(\mathsf{E}f)(x_{t})$ is the expected value of $f$ one step ahead in the future, given that at time $t$ the system was found in $x_{t}$
$(\mathsf{E}f)(x_{t})=\int p(dy|x_{t})f(y)=\mathbb{E}_{y\sim X_{t+1}|X_{t}}[f(y)|x_{t}].$
Notice that $\mathsf{E}$ is an operator because it maps any function $f$ to another function, $x_{t}\mapsto(\mathsf{E}f)(x_{t})$, and is linear because $\mathsf{E}(f+\alpha g)=\mathsf{E}f+\alpha\mathsf{E}g$. When the dynamics is deterministic, $\mathsf{E}$ is known as the Koopman operator (Koopman, 1931), while in the stochastic case it is known as the transfer operator (Applebaum, 2009). Arguably, the most important feature of evolution operators is their spectral decomposition (Mezić, 2005), which can be used to express the dynamics as a linear superposition of modes. These ideas lie at the core of the celebrated Time-lagged Independent Component Analysis (Molgedey and Schuster, 1994), and Dynamical Mode Decomposition (DMD) (Schimid, 2010; Kutz et al., 2016).
Evolution operator learning is best understood from the perspective of latent linear dynamical models, which is schematically depicted in Figure 1. In this framework, the dynamical state $x_{t}$ is first mapped into a latent space defined by a (fixed or learned) representation $\varphi$. Then, a linear evolution map $E$ is learned to approximate the dynamics of the latents. The pair $(\varphi,E)$ provides an approximation of $\mathsf{E}$ restricted to the $d$-dimensional subspace spanned by the components of $\varphi$, given the data. kooplearn implements state-of-the-art methods to learn $\varphi$, $E$, and the associated spectral decomposition of $\mathsf{E}$.
Figure 1: Sketch of the action of an evolution operator on a protein folding trajectory. The dynamics of the protein is linearized by means of a nonlinear representation $\varphi$ and consequently evolved by means of the linear map $E$.
The ecosystem of Python libraries that support operator-based modeling has grown considerably in recent years, with a predominant focus on the DMD family of methods. PyDMD (Ichinaga et al., 2024) emphasizes classical and kernel DMD variants; pykoopman (Pan et al., 2024) implements classical DMD methods with dictionary-based feature maps; pykoop (Dahdah and Forbes, 2025) offers a modular framework for lifting-function construction with a focus on system identification and control; DLKoopman (Dey and Davis, 2023) focuses on autoencoder approaches, while KoopmanLab (Xiong et al., 2023) targets Koopman neural operators. kooplearn addresses the general problem of learning evolution operators, and it is the result of a multi-year research effort in innovative operator learning algorithms. While it provides standard prediction and spectral decomposition utilities, it extends the state of the art in evolution operator learning codes by implementing fast kernel estimators (Meanti et al., 2023; Turri et al., 2025b), infinitesimal generator models for SDEs (Kostic et al., 2024d; Devergne et al., 2024), and specialized losses for deep representation learning (Mardt et al., 2018; Kostic et al., 2024c, a; Turri et al., 2025a). We now provide a concise overview of the functionality of kooplearn.
Learning Linear Evolution Maps $E$
kooplearn implements state-of-the-art algorithms for learning evolution operators when the representation $\varphi$ is fixed. The library offers estimators in both their linear and kernel formulations (see the Ridge and KernelRidge classes), which bridge the gap between recent theoretical advances (Kostic et al., 2022, 2023, 2024b, 2024c) and practical code implementations. A key model in kooplearn is the kernel-based Reduced Rank Regression (Kostic et al., 2022). This estimator provably outperforms traditional methods (Williams et al., 2015) in approximating the operator’s spectrum (Kostic et al., 2023), as illustrated in Figure 2. To our knowledge, kooplearn provides the only open-source implementation of this algorithm. To handle large datasets, kooplearn also includes randomized (Turri et al., 2025b) and Nyström-based (Meanti et al., 2023) kernel estimators, which significantly speed up the fitting process, making it one of the fastest libraries for kernel-based operator learning, as shown in Figure 3.
Figure 2: Comparison between kernel DMD (kDMD) and Reduced Rank estimators. The Reduced Rank estimator provides a more accurate approximation of the leading eigenfunctions of the transfer operator for the overdamped Langevin dynamics.
Figure 3: Fit time of a Kernel model (Gaussian kernel) on a dataset of 50005000 observations from the Lorenz 63 dynamical system. The results are the median of three independent runs on a system equipped with an Intel Core i9-9900X CPU (3.50GHz) and 48GB of RAM memory.
Learning the Representation $\varphi$
kooplearn also exposes theoretically-grounded loss functions — implemented in both PyTorch (Paszke et al., 2019) and JAX (Bradbury et al., 2018) — suited for learning the representation $\varphi$ with neural network models. This allows the incorporation of structural priors, such as graph-based encoders. Within this deep-learning approach, two main families are supported: (i) encoder-decoder schemes with the loss proposed in Lusch et al. (2018), and (ii) encoder-only schemes, for which kooplearn implements the VAMP loss (Mardt et al., 2018) and the spectral contrastive loss (Turri et al., 2025a).
Learning the Infinitesimal Generator of Diffusion Processes
In continuous-time dynamics, the system’s evolution operator can be expressed as the exponential of the infinitesimal generator $\mathsf{L}$, a differential operator defined by the equations of motion (Applebaum, 2009, Chapter 3). Formally, for time-homogeneous dynamics, the generator relates to the evolution operator via $\mathsf{E}=e^{\mathsf{L}}$, and consequently $\mathbb{E}[f(X_{t})|x_{0}]=(e^{t\mathsf{L}}f)(x_{0})$. Since the exponential of an operator preserves its eigenfunctions, one can use knowledge of $\mathsf{L}$ (or its properties) to learn dynamical behavior without requiring lag-time data. In other words, it becomes possible to construct a physics-informed kinetic model $\mathsf{E}$ solely from static (equilibrium) data. To this end, kooplearn provides implementations of recent kernel-based algorithms for diffusion processes with Dirichlet boundary conditions from Kostic et al. (2024d), as well as neural representations as proposed in Devergne et al. (2024). As demonstrated in Devergne et al. (2025), these approaches improve sample complexity compared to estimators that rely solely on lag-time trajectory data.
Datasets
Figure 4: Samples from the datasets included in kooplearn.
To foster reproducibility and rigorous benchmarking, kooplearn includes the kooplearn.datasets module, containing utilities to easily generate trajectories for systems that range from deterministic chaos (e.g., Lorenz-63, Duffing oscillator, Logistic Map) to stochastic and metastable dynamics (e.g., stochastic linear systems, regime-switching models, Langevin dynamics). A distinguishing feature of the library is the inclusion of benchmarks with accessible ground-truth spectral decompositions—such as the Noisy Logistic Map (Ostruszka et al., 2000) and Overdamped Langevin Dynamics in a quadruple-well potential (Prinz et al., 2011). These allow users to quantify the accuracy of learned eigenvalues and eigenfunctions directly (as demonstrated in Figure 2). Finally, the suite includes the Ordered MNIST from (Kostic et al., 2022) to evaluate performance on high-dimensional structured data. Examples of trajectories generated using the kooplearn.datasets module are illustrated in Figure 4.
Conclusion
kooplearn closely follows the scikit-learn API (Pedregosa et al., 2011) and strives to lower the technical barrier to experimenting with evolution operators. At the same time, it provides optimized implementations of state-of-the-art algorithms for evolution operator learning, making it valuable for research, education, rapid prototyping, and exploratory analysis of dynamical systems. As of today, kooplearn has been employed in a variety of studies (Kostic et al., 2022; Bevanda et al., 2023; Kostic et al., 2023, 2024b, 2024c; Turri et al., 2025a; Bevanda et al., 2025). It can be installed using the command pip install kooplearn. Its documentation, alongside many worked-out examples, is available on the webpage https://kooplearn.readthedocs.io/.
Acknowledgements
This work was partially funded by the European Union - NextGenerationEU and by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), through the PNRR MUR Project PE000013 CUP J53C22003010006 "Future Artificial Intelligence Research (FAIR)" and EU Project ELIAS under grant agreement No. 101120237.